OpenStudy (samigupta8):

If the displacement of a particle varies with time as √x= t + 7, then velocity of particle is A) inversely proportional to x B) proportional to x C) proportional to √x D) inversely proportional to √x

OpenStudy (anonymous):

It is immensely easy to solve. take a partial derivative from both sides...

OpenStudy (samigupta8):

wt then...???

OpenStudy (anonymous):

Wait a minute ...

OpenStudy (anonymous):

\[\sqrt{x}=t+7 \rightarrow x=(t+7)^2 \rightarrow v=\frac{ dx }{ dt }=2(t+7) \rightarrow v=2\sqrt{x}\]

OpenStudy (samigupta8):

pkk.....i got it....

OpenStudy (anonymous):

Taking partial derivatives of both sides gives =d/dt(x^1/2)=d/dt(t+7) =d/dx(x^1/2)dx/dt=1 =1/2x^(-1/2)dx/dt=1 -->dx/dt=2x^1/2 so C can be true...